mhchen:

Linear Algebra: Define the fibonnaci sequence as such: F(0) = 0 F(1) = 1 for all n > 1, F(n) = F(n-1) + F(n-2) Prove this statement with induction: Using proof by induction:

mhchen:

If \[Q = \left[\begin{matrix}F_{2} & F_{1} \\ F_{1} & F_{0}\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]\] Then \[Q^{n} = \left[\begin{matrix}F_{n+1} & F_{n} \\ F_{n} & F_{0}\end{matrix}\right]\]

mhchen:

Let \[P(n) = Q^{n}\] Then \[P(1) = Q^{1} = \left[\begin{matrix}F_{1+1} & F_{1} \\ F_{1} & F_{0}\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right] \] And assuming P(n) is true, \[P(n+1) = Q^{n+1} = Q^{n}Q^{1} = \left[\begin{matrix}F_{n+1} & F_{n} \\ F_{n} & F_{0}\end{matrix}\right]\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right] = \left[\begin{matrix}F_{n+1}+F_{n} & F_{n+1} \\ F_{n+1}+F_{0} & F_{n}\end{matrix}\right] \] \[= \left[\begin{matrix}F_{n+2} & F_{n+1} \\ F_{n+1} & F_{n}\end{matrix}\right]\] but from the last equation, the bottom-right entry should be F_{0} not F_{n} so I don't know what happened.

akshay:

@Narad He can help in Linear Algebra

mhchen:

Looking at this: https://math.stackexchange.com/questions/867394/how-to-compute-the-nth-number-of-a-general-fibonacci-sequence-with-matrix-multip I feel like the textbook I'm using is wrong, and I'm right xd

akshay:

Hey man, after all, textbooks are written by humans and our intellect rules over that textbook. Who knows, maybe it is wrong! xD

mhchen:

Based on this website, I'm also right and the textbook is wrong: https://www.slader.com/textbook/9781118473504-elementary-linear-algebra-11th-edition/51/technology-exercises/3/ I'm going to contact my classmates to see if they noticed.

mhchen:

Wait no I messed up on the bottom-left entry

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mhchen And assuming P(n) is true, \[P(n+1) = Q^{n+1} = Q^{n}Q^{1} = \left[\begin{matrix}F_{n+1} & F_{n} \\ F_{n} & F_{0}\end{matrix}\right]\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right] = \left[\begin{matrix}F_{n+1}+F_{n} & F_{n+1} \\ F_{n}+F_{0} & F_{n}\end{matrix}\right] \] \[= \left[\begin{matrix}F_{n+2} & F_{n+1} \\ F_{n} & F_{n}\end{matrix}\right]\] \(\color{#0cbb34}{\text{End of Quote}}\)

Gdeinward:

I hate matrixes

mhchen:

The teacher confirmed, textbook was wrong

Gdeinward:

you gotta know your smart, when you find an error with the textbook